Hydrogen Cyanide Poisoning Part A Calculate the amount of HCN that gives the lethal dose in a small laboratory room measuring 12.0 ftx 15.0 ft × 960ft. Hydrogen cyanide, HCN, is a poisonous gas. The ethal dose is approximately 300. mg kilogram of air when inhaled. The density of air at 26。C is 0.00118 g/cm per Express your answer to three significant figures and include the appropriate units. Hints Value Units
Answer #1

we will use below conversion.
1 ft = 30.48 cm

volume = 12.0 ft * 15.0 ft * 9.60
= 1728 ft^3
= 1728 * (30.48 cm)^3
= 1728 * 28317 cm^3
= 4.893*10^7 cm^3

now use:
density of air = mass of air / volume
0.00118 g/cm^3 = mass of air / (4.893*10^7 cm^3)
mass of air = 5.774*10^4 g
mass of air = 57.74 Kg
In 1 Kg of air, mass of HCN allowed = 300 mg
mass of HCN allowed = 57.74*300 mg = 17322 mg = 17.3 g
Answer: 17.3 g